### v1

@@ Abstract
Agglomeration techniques are important to reduce the computational costs of numerical simulations and stand at the basis of multilevel algebraic solvers. To automatically perform the agglomeration of polyhedral grids, we propose a novel Geometrical Deep Learning-based algorithm that can exploit the geometrical and physical information of the underlying computational domain to construct the agglomerated grid and simultaneously guarantee the agglomerated grid’s quality. In particular, we propose a bisection model based on Graph Neural Networks (GNNs) to partition a suitable connectivity graph of computational three-dimensional meshes. The new approach has a high online inference speed and can simultaneously process the graph structure of the mesh, the geometrical information of the mesh (e.g. elements’ volumes, centers’ coordinates), and the physical information of the domain (e.g. physical parameters). Taking advantage of this new approach, our algorithm can agglomerate meshes of a domain composed of heterogeneous media in an automatic way. The proposed GNN techniques are compared with the k-means algorithm and METIS: standard approaches for graph partitioning that are meant to process only the connectivity information on the mesh. We demonstrate that the performance of our algorithms outperforms available approaches in terms of quality metrics and runtimes. Moreover, we demonstrate that our algorithm also shows a good level of generalization when applied to more complex geometries, such as three-dimensional geometries reconstructed from medical images. Finally, the capabilities of the model in performing agglomeration of heterogeneous domains are tested in the framework of problems containing microstructures and on a complex geometry such as the human brain.
@@ 1 Introduction
Nowadays, the numerical solution of differential problems is central to all fields of engineering and applied sciences, moreover, it is expanding its boundaries to other fields which can benefit from numerical simulations. The applications in all these fields often involve complex physical domains, possibly featuring moving geometries and interfaces, heterogeneous media, immersed interfaces [1], and complex structures [2, 3]. The need to deal with such complex structures leads to the development of new and more sophisticated Finite Element Methods (FEM) that can employ general polygons and polyhedra as grid elements for the numerical discretization of Partial Differential Equations (PDEs). We mention the Polyhedral Discontinuous Galerkin method [4, 5, 6, 7, 8, 9], the Virtual Element Method [10, 11, 12, 13, 14], the Hybrid High-Order method [15, 16, 17, 18], the mimetic finite difference method [19, 20, 21, 22], and the hybridizable discontinuous Galerkin method [23, 24, 25, 26].
@@
The flourishing of these methods takes with it the possibility of an intensive usage of mesh agglomeration techniques, that allow to merge mesh elements to obtain coarser polytopal grids. However, the issue of efficiently performing agglomeration of polytopal meshes is still an open problem and object of intense research [27, 28]. Mesh agglomeration is a powerful tool because it can reduce the mesh computational complexity while retaining a detailed geometrical description of the computational domain and/or interfaces. In scientific simulations, reducing the number of elements of a computational mesh can lead to faster simulations and quicker results while maintaining accuracy for example locally increasing the polynomial order. Furthermore, agglomerated meshes stand at the basis of generating a hierarchy of (nested) coarser grids starting from a fine mesh of a complex physical domain of interest. Then such grids’ sequence, characterized by different refinement levels, can be employed in multi-grid solvers [29, 30, 31, 32, 33, 34, 35] to accelerate the numerical solution of the underlying algebraic system.
@@
Because of the impossibility of the agglomerated grids to be used in the framework of classical FEMs, grid agglomeration is a quite unexplored topic. Currently employed agglomeration methods rely on iterative algorithms based on graph partitioning of the connectivity matrix, which make them not particularly feasible for complicated or heterogeneous domains or whenever the underlying PDE models multiple physical processes, that have to be properly taken into account in the agglomeration process. Moreover, during the agglomeration procedure, the quality preservation of the underlying mesh is fundamental because it might affect the overall stability and accuracy of the numeric method. Indeed, a suitably agglomerated mesh can give the same accuracy of the solution computed on the finer grid, but with less degrees of freedom. This leads to memory savings and lower computational loads.
@@
In recent years the use of Machine Learning (ML) techniques to accelerate numerical methods has become widespread [36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. In this work, we present a novel geometric deep-learning approach to address the problem of agglomerating three-dimensional meshes of heterogeneous domains. The proposed approach is based on Graph Neural Networks (GNNs). Indeed, we do not simply try to decide an a-priori criterium to agglomerate the mesh, which would inevitably result in poor performance or high computational cost due to the impossibility of capturing all the possible configurations. Instead, our ML strategy exploits and processes automatically the huge amount of available data to learn only the distribution of the features of interest for the application, leading to high performance and computational efficiency.
@@
In our approach, the problem of mesh agglomeration is re-framed as a graph partitioning problem by exploiting the structure of the mesh. More precisely, the graph representation of the mesh is obtained by assigning a node to each mesh element. Edges connect pairs of nodes relative to adjacent mesh elements (see Figure 1). By exploiting such an equivalent representation, we can apply the proposed ML algorithm to solve a node classification problem, where each element is assigned to a cluster, corresponding to an element of the agglomerated mesh.
@@
Furthermore, a powerful ability of the proposed GNNs-based algorithm is the possibility to consider further physical properties of the computational domain during mesh agglomeration. Indeed, by adding new input features, and slightly modifying the network structure (particularly, changing the loss function), we can construct an algorithm that takes into account the heterogeneities of the physical domain or possible physical features (such as e.g. embedded microstructures). By exploiting this idea, we can develop a model that automatically agglomerates heterogeneous domains, without the need to separately agglomerate regions with different material properties, with the need for a post-process to glue the obtained meshes to create the final one. Moreover, this new automatic algorithm can have an even bigger impact if the computational domain exhibits, for example, micro-structures and thus cannot be divided into simpler sub-regions.
In this work, we demonstrate that the proposed GNNs-based agglomeration algorithm has high online inference speed and it can process naturally and simultaneously both the graph structure of mesh and the geometrical information that can be attached to the nodes, such as the elements areas/volumes or their centers’ coordinates. The technique is then compared with $k$-means [46] algorithm, with another ML algorithm that can process only the geometrical information, and with METIS [47], a standard approach for graph partitioning that is meant to account only for the graph information on the mesh. The study extends the results of [27], which were not directly extended for agglomerating three-dimensional meshes or domains with complicated structures.
Finally, we remark that the proposed algorithm is general and the training can be done once and for all. More precisely, the models do not depend either on the PDE under investigation nor on the numerical method adopted for the discretization.
@@
The rest of the paper is organized as follows. In Section 2, we present the GNN-agglomeration focusing on homogeneous domains in Section 2.1, and on heterogeneous ones in Section 2.2. In Section 3, we present the numerical results focusing on homogeneous domains, first by measuring the effectiveness of the proposed strategies in terms of computational cost, and quality metrics, then evaluating the generalization capabilities over a complex domain such as the human brain ventricles. In Section 4, we focus on the numerical results of heterogeneous domains, by testing the quality of simple datasets and complex heterogeneous domains: one domain containing microstructures and a human brain. Finally, in Section 5 we draw some conclusions.
@@ 2 Mesh agglomeration algorithms via GNN
In this section, we introduce our agglomeration algorithm for three-dimensional meshes, discussing also the treatment of heterogeneous media. Most of agglomeration algorithms tackle the problem by re-framing it into a graph partitioning problem by exploiting the connectivity structure of the mesh. In this work, we will focus on constructing an agglomeration algorithm, based on recursive bisection of the associated graph. This strategy recursively bisects the input mesh’s connectivity graph until the agglomerated elements have the desired size, as explained in Algorithm 1.
@@
     $Y\leftarrow\mathscr{M}(G,X)$
@@
In the particular case of our strategy, we propose to employ GNN-based bisection models $\mathscr{M}$ that take as input a graph $G$ together with a set of features $X$ attached to each node (e.g. barycentric coordinates, area/volume and eventually physical parameters) and output the vector of probabilities $Y$ of each node to belong to cluster $1$ or $2$. This strategy has key advantages with respect to the state-of-the-art methods (METIS [47] and k-means [46]), more precisely:
@@
  • It can process information about the connectivity graph of mesh elements, geometrical features of the elements, and physical properties of the modeled material. On the contrary, METIS is meant to process only graph-based information, and k-means only geometrical information.
@@
  • The online inference on new instances is much faster than METIS and k-means. Indeed, in our case, the main computational burden is faced during the training phase, while the other algorithms perform iterations online to provide an estimate of the solution. Moreover, the training phase is independent of the geometry and of the underlying PDE model, so the training phase is performed once and for all.
@@
  • The GNN automatically processes additional information without requiring the user to know how that information can benefit performance. On the contrary, for METIS and k-means this has to be explicitly modeled.
@@
Notice that GNN-drive agglomeration techniques were proposed in [27], for a two-dimensional setting and homogeneous domain. However, the approach in [27] was limited to two-dimensional cases. The main novelty of the extension considered here is the possibility to handle also three-dimensional settings and properly take into account heterogeneities and/or embedded interfaces. More precisely, to handle the variability of the 3D meshes, which are much more complex than the 2D ones, we propose suitable improvements with respect to the algorithm proposed in [27]. Those improvements are both related to the pipeline and the GNN’s architecture. Moreover, when considering a three-dimensional setting, training and inference need to be feasible with much larger datasets, due to the broader amount of elements that 3D meshes contain, even just in simple geometries. Moreover, we propose also a GNN architecture which is able to treat the agglomeration of heterogeneous domains, changing the loss function of the network.
In Figure 2 we report a schematic graphical conceptualization of the workflow of the proposed GNN agglomeration algorithm. In all cases the dataset of meshes is created by means of GMSH [48]. Then, using a Python code, the graph is extracted for each mesh and the dataset is created by bundling graphs adjacencies and features matrices) together. Then the GNN model is trained in an unsupervised learning setting, using a PyTorch Geometric [49] implementation. Finally, the trained model is loaded in evaluation mode and used to perform the graph bisection iteratively, obtaining a clustered graph, which is then post-processed in a Matlab environment to build the agglomerated mesh.
Figure 2: Schematic representation of the agglomeration pipeline
@@ 2.1 Architectures of GNN-agglomeration models for homogeneous domains
In this section, we present the proposed architecture of the GNN for mesh agglomeration for homogeneous domains. In particular, we explain the proposed architecture (GNN-AGGL-Enhanced), comparing it with the simplest possible extension of the one proposed in [27] (GNN-AGGL-Base). Indeed, graph partitioning is a problem that focuses on the structure of the graph itself regardless of any specific spatial representation. This implies that for graphs in two or three dimensions, the principles and algorithms for graph partitioning remain the same, while obviously, suitable modifications to handle 3D data are needed. However, from the comparison of the result reported in Section 3 below, it is evident the need of a more complex GNN architecture, as the GNN-AGGL-Enhanced one, for the agglomeration task in 3D.
The proposed models (GNN-AGGL-Base and GNN-AGGL-Enhanced) are both based on two GNN SAGE-Base models, extending the ideas of [27], to the three-dimensional space. Both graph-bisection models perform a classification task taking as input the graph $G=(V,E)$ and the features related to its nodes $X\in\mathbb{R}^{N\times 4}$. The columns of the input feature matrix $X$ at the $i$-th row correspond to the information of the $i$-th mesh element. In particular, we have:
@@
where $(x_{i},y_{i},z_{i})$ are the $(x,y,z)$-coordinates of the $i$-th element’s centroid, while the $V_{i}$ is the element’s volume. The output is $Y\in\mathbb{R}^{N\times 2}$, where the first column represents the probability of belonging to the class $S_{1}$, while the second is the probability of being in the class $S_{2}$ of the partition. The characteristics of the resulting networks are reported in Table 1. We underline that the number of parameters of the GNN-AGGL-Enhanced model is $\sim 110\mathrm{K}$, with respect to the $\sim 28\mathrm{K}$ of the GNN-AGGL-Base model. However, as will be depicted in Section 3, this level of complexity is required to correctly tame the inherited geometric complexity.
@@
 •  Coordinates rescaled in $[-1,1]$.  •  Coordinates standardization:
Normalization        Mean: $\mu=0$ Variance: $\sigma^{2}=1$
@@
Linear Layer 3 decreasing-layers of $32-8-2$ neurons. 4 decreasing-layers of $64-32-8-2$ neurons.
@@
The training of the models has been performed in an unsupervised learning setting. The loss function that is used is the expected normalized cut. For details regarding the loss function construction, we refer to [27].
Concerning the construction of the dataset, we generate a training dataset of $500$ tetrahedral meshes ($125$ meshes of the unitary cube and $375$ meshes of suitable portions of them). All the meshes are generated using GMSH [48]. We adopt a split between the training set and validation set of respectively the $80\%$ and the $20\%$ of the total dataset. As a consequence, The training set is composed of $400$ grids ($100$ meshes of the unitary cube and $300$ meshes of portions), while the validation dataset is made by the remaining $100$ grids. To guarantee good agglomeration results with our GNN models, we use a dataset with a large variety of meshes, differentiating the number of mesh elements. In the training and validation datasets, we start from coarse meshes of $600$ elements, until around 42 0004200042\,00042 000 elements of mesh in the finest one. In Figure 3, we report some examples of meshes from the training set.
@@
  • Regularization, summing an $L^{2}$-regularization term to the normalized cut loss, leading to the following loss definition:
    ^$\hat{\mathscr{L}}(W,Y,G)=\mathscr{L}(W,Y,G)+\lambda\|W\|^{2}_{2}.$
@@
The training process has been performed by adopting an Adam optimizer. All the details about the training hyper-parameters for the two models are reported in Table 2.
@@
Batch size $4$ Samples $4$ Samples
@@
Table 2: Training hyper-parameters adopted for the two GNN proposed models.
0050505050100100100100150150150150200200200200250250250250300300300300001111222233334444Epochs^$\hat{\mathscr{L}}(W,Y,G)$Loss function for the GNN-AGGL-Base modelTraining lossValidation loss
0050505050100100100100150150150150200200200200250250250250300300300300350350350350400400400400001111222233334444Epochs^$\hat{\mathscr{L}}(W,Y,G)$Loss function for the GNN-AGGL-Enhanced modelTraining lossValidation loss
@@
In Figure 4, we report the values of the loss functions for the training and validation sets for the GNN-AGGL-Base and the GNN-AGGL-Enhanced models. We can observe that using the GNN-AGGL-Enhanced model, the loss function reaches lower values on both datasets. Moreover, the training procedure is more stable in contrast to what happens in the two-dimensional setting [27]. This is probably a consequence of the fact that the grids of the three-dimensional training dataset are made only of tetrahedra and not of general polyhedra.

### v2

@@ Abstract
Agglomeration techniques can be successfully employed to reduce the computational costs of numerical simulations and stand at the basis of multilevel algebraic solvers. To automatically perform mesh agglomeration, we propose a novel Geometrical Deep Learning-based algorithm that can exploit the geometrical and physical information of the underlying computational domain to construct the agglomerated grid and -simultaneously-guarantee the agglomerated grid’s quality. In particular, we propose a bisection model based on Graph Neural Networks (GNNs) to partition a suitable connectivity graph of computational three-dimensional meshes. The new approach has a high online inference speed. It can simultaneously process the graph structure of the mesh, the geometrical information of the mesh (e.g., elements’ volumes, centers’ coordinates), and the physical information of the domain (e.g., physical parameters). Taking advantage of this new approach, our algorithm can agglomerate meshes of a domain composed of heterogeneous media, automatically respecting the underlying heterogeneities. The proposed GNN approach is compared with the k-means algorithm and METIS, which are widely employed approaches for graph partitioning and are meant to process only the connectivity information on the mesh. We demonstrate that the performance of our algorithms outperforms the k-means and METIS algorithms in terms of quality metrics and runtimes. Moreover, we demonstrate that our algorithm also shows a good level of generalization when applied to complex geometries, such as three-dimensional geometries reconstructed from medical images. Finally, the model’s capability to perform agglomeration in heterogeneous domains is evaluated when integrated into a polytopal discontinuous Galerkin finite element solver.
@@ 1 Introduction
Nowadays, the numerical solution of differential problems is central to all fields of engineering and applied sciences; moreover, it is expanding its boundaries to other fields that can benefit from numerical simulations. The applications in all these fields often involve complex physical domains, possibly featuring moving geometries and interfaces, heterogeneous media, immersed interfaces [1], and complex structures [2, 3]. The need to deal with such complex structures leads to the development of new and more sophisticated Finite Element Methods (FEM) that can employ general polygons and polyhedra (polytopal, in short) as grid elements for the numerical discretization of Partial Differential Equations (PDEs). We mention the Polyhedral Discontinuous Galerkin method [4, 5, 6, 7, 8, 9], the Virtual Element Method [10, 11, 12, 13, 14], the Hybrid High-Order method [15, 16, 17, 18], the mimetic finite difference method [19, 20, 21, 22], and the hybridizable discontinuous Galerkin method [23, 24, 25, 26]. The flourishing of these methods takes with it the possibility of an intensive use of mesh agglomeration techniques that allow merging mesh elements to obtain coarser polytopal grids. However, the issue of efficiently performing agglomeration of polytopal meshes is still an open problem and object of intense research [27, 28]. Mesh agglomeration is a powerful tool because it can reduce the computational complexity of the mesh while retaining a detailed geometrical description of the computational domain and/or interfaces. In scientific simulations, reducing the number of elements of a computational mesh can lead to faster simulations and quicker results while maintaining accuracy, for example, by locally increasing the polynomial order. Furthermore, agglomerated meshes form the basis for generating a hierarchy of (nested) coarser grids starting from a fine mesh of a complex physical domain of interest. Then, such grids’ sequence, characterized by different refinement levels, can be employed in multi-grid solvers [29, 30, 31, 32, 33, 34, 35] to accelerate the numerical solution of the underlying algebraic system.
@@
Because of the impossibility of the agglomerated grids to be used in the framework of classical FEMs, grid agglomeration is a relatively unexplored topic. Currently employed agglomeration methods rely on iterative algorithms based on graph partitioning of the connectivity matrix, which makes them not particularly feasible for complicated or heterogeneous domains or whenever the differential system models multiple physical processes that have to be appropriately taken into account in the agglomeration process. Moreover, during the agglomeration procedure, the quality preservation of the underlying mesh is fundamental because it might affect the overall stability and accuracy of the numerical method. Indeed, a suitably agglomerated mesh can give the same accuracy of the solution computed on the finer grid, but with fewer degrees of freedom. This leads to memory savings and lower computational loads.
@@
In recent years the use of Machine Learning (ML) techniques to accelerate numerical methods has become widespread [36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. In this work, we present a novel geometric deep-learning approach to address the problem of agglomerating three-dimensional meshes of heterogeneous domains. The proposed approach is based on Graph Neural Networks (GNNs). Indeed, we do not simply try to decide an a-priori criterion to agglomerate the mesh, which would inevitably result in poor performance or high computational cost due to the impossibility of capturing all the possible configurations. Instead, our ML strategy automatically exploits and processes the huge amount of available data to learn only the distribution of the features of interest for the application, leading to high performance and computational efficiency. Moreover, the strong generalization capabilities of GNN architectures in partitioning enable the agglomeration of large meshes using only small meshes in the training set, thereby significantly reducing training costs, as discussed in [46].
@@
In our approach, the problem of mesh agglomeration is re-framed as a graph partitioning problem by exploiting the mesh structure. More precisely, the graph representation of the (dual) mesh is obtained by assigning a node to each mesh element. Edges connect pairs of nodes relative to adjacent mesh elements (see Figure 1). By exploiting such an equivalent representation, we can apply the proposed ML algorithm to solve a node classification problem, where each element is assigned to a cluster corresponding to a new polyhedral cell in the agglomerated mesh. Furthermore, a powerful ability of the proposed GNN-based algorithm is the possibility to consider further physical properties of the computational domain during the process of mesh agglomeration. Indeed, by adding suitable input features and slightly modifying the network structure (particularly, changing the loss function), we can construct an algorithm that takes into account the heterogeneities of the underlying physical domain or possible physical features (e.g., embedded layers and/or microstructures). By exploiting this idea, we can develop a model that automatically agglomerates heterogeneous domains without the need to separately agglomerate regions with different material properties, with the need for a post-process to glue the obtained meshes to create the final one. Moreover, this new automatic algorithm can have an even bigger impact if the computational domain exhibits, for example, micro-structures and thus cannot be divided into simpler sub-regions. Further, we demonstrate that the proposed GNNs-based agglomeration algorithm has high online inference speed and can process naturally and simultaneously both the graph structure of the mesh and the geometrical information that can be associated with the graph nodes, such as the elements’ areas/volumes or their centers’ coordinates. The proposed GNNs-based agglomeration algorithm is then compared with the $k$-means [47] algorithm (another ML algorithm that can process only the geometrical information), and with METIS [48], a standard approach for graph partitioning that is meant to account only for the graph information on the mesh. The study extends the results of [27], which cannot be directly extended for agglomerating three-dimensional meshes or domains with complicated structures. Finally, we remark that the proposed algorithm is general, and the training can be done once and for all. More precisely, the models do not depend on the PDE under investigation or the numerical method adopted for the discretization.
@@
The rest of the paper is organized as follows. In Section 2, we present the GNN-agglomeration focusing on homogeneous domains in Section 2.1, and on heterogeneous ones in Section 2.2. In Section 3, we present the numerical results focusing on homogeneous domains, first by measuring the effectiveness of the proposed strategies in terms of computational cost and quality metrics, then evaluating the generalization capabilities over a complex domain such as the human brain ventricles. Section 4 focuses on the numerical results of heterogeneous domains by testing the quality of simple datasets and complex heterogeneous domains. Section 6 presents numerical results demonstrating the performance of the proposed agglomeration strategy when the resulting meshes are used within PDE solvers. Finally, in Section 7, we draw some conclusions.
@@ 2 Mesh agglomeration algorithms via GNN
In this section, we introduce our agglomeration algorithm for three-dimensional meshes, discussing also the treatment of heterogeneous media. Most of the agglomeration algorithms tackle the problem by re-framing it into a graph partitioning problem by exploiting the connectivity structure of the mesh. In this work, we will focus on constructing an agglomeration algorithm based on recursive bisection of the associated graph. This strategy recursively bisects the input mesh’s connectivity graph until the agglomerated elements have the desired size, as explained in Algorithm 1.
@@
  Obtain the partition $Y$ by means of the bisection model $\mathscr{M}(G,X)$ (i.e., GNN, METIS or k-means).
@@
In the particular case of our strategy, we propose to employ GNN-based bisection models $\mathscr{M}$ that take as input a graph $G$ together with a set of features $X$ attached to each node (e.g. barycentric coordinates, area/volume and eventually physical parameters) and output the vector of probabilities $Y$ of each node to belong to cluster $1$ or $2$. This strategy has key advantages compared to some of the state-of-the-art methods like METIS [48] and k-means [47]. More precisely:
@@
  • It can process information about the connectivity graph of mesh elements, geometrical features of the elements, and physical properties of the modeled material. On the contrary, METIS is meant to process only graph-based information, and considering the geometrical and physical parameters would require modification of graph weights, which could not be straightforward to define in real applications. The k-means algorithm is typically used to process only geometrical information and does not take into account the connectivity matrix of the mesh.
@@
  • The online inference on new instances is much faster than METIS and k-means. Indeed, in our case, the main computational burden is faced during the training phase, while the other algorithms perform iterations online to provide an estimate of the solution. Moreover, the training phase is independent of the geometry and the underlying PDE model, so it is performed once and for all.
@@
  • The GNN automatically processes additional information without requiring the user to know how that information can benefit performance. On the contrary, for METIS and k-means, this has to be explicitly modeled.
@@
We point out that in this study, we restrict our comparison to the k-means and METIS algorithms. A more comprehensive benchmarking (in terms of robustness and scalability) with other advanced, domain-specific partitioning strategies such as Coarsening Schemes for Graph Partitioning (e.g. [49, 50]), or SCOTCH [51, 52], which are known for their effectiveness in specific scenarios, falls beyond the scope of this work but represents an important objective for future research. Notice that GNN-driven agglomeration techniques were first proposed in [27], for a two-dimensional setting and underlying homogeneous domains; however, the approach in [27] was limited to two-dimensional cases. The main novelty of the extension considered here is the possibility to handle also three-dimensional settings and properly take into account heterogeneities and/or embedded interfaces. More precisely, to handle the variability of three-dimensional meshes, which are in general much more complex, we propose suitable improvements with respect to the algorithm proposed in [27]. Those improvements are both related to the pipeline and the GNN’s architecture. Moreover, when considering a three-dimensional setting, training and inference need to be feasible with much larger datasets, due to the broader amount of elements that three-dimensional meshes contain, even just in simple geometries. Moreover, we also propose a GNN architecture that can treat the agglomeration of heterogeneous domains, changing the loss function of the network. In Figure 2, we report a schematic graphical conceptualization of the workflow of the proposed GNN agglomeration algorithm. In all cases, the dataset of meshes is created using GMSH [53]. Then, the graph is extracted for each mesh, and the dataset is created by bundling graph adjacencies and feature matrices together. Next, the GNN model is trained in an unsupervised learning setting using a PyTorch Geometric [54] implementation. Finally, the trained model is loaded in evaluation mode and used to perform the graph bisection iteratively, obtaining a clustered graph, which is then post-processed to build the agglomerated mesh.
Figure 2: Schematic representation of the agglomeration pipeline.
@@ 2.1 Architectures of GNN-agglomeration models for homogeneous domains
In this section, we present the proposed GNN architecture for mesh agglomeration for homogeneous domains. In particular, we describe the proposed architecture (GNN-AGGL-Enhanced), comparing it with the most straightforward possible three-dimensional extension of the one proposed in [27] (GNN-AGGL-Base). Indeed, graph partitioning is a problem that focuses on the structure of the graph itself, regardless of any specific spatial representation. This implies that for graphs in two or three dimensions, the principles and algorithms for graph partitioning remain the same. At the same time, suitable modifications to handle three-dimensional data are needed. However, from the comparison of the result reported in Section 3 below, it is evident that there is a need for a more complex GNN architecture, as the GNN-AGGL-Enhanced one, for the agglomeration task in three dimensions.
The proposed models (GNN-AGGL-Base and GNN-AGGL-Enhanced) are both based on two GNN SAGE-Base models, extending the ideas of [27] to three-dimensional space. Both graph-bisection models perform a classification task, taking as input the graph $G=(V,E)$ and the features related to its nodes $X\in\mathbb{R}^{N\times 4}$. The columns of the input feature matrix $X$ at the $i$-th row correspond to the information of the $i$-th mesh element. In particular, we have:
@@
where $(x_{i},y_{i},z_{i})$ are the $(x,y,z)$-coordinates of the $i$-th element’s centroid, while the $V_{i}$ is the element’s volume. The output is $Y\in\mathbb{R}^{N\times 2}$, where the first column represents the probability of belonging to the class $S_{1}$, while the second is the probability of being in the class $S_{2}$ of the partition. The characteristics of the resulting networks are reported in Table 1. We underline that the number of parameters of the GNN-AGGL-Enhanced model is $\sim 110\mathrm{K}$, with respect to the $\sim 28\mathrm{K}$ of the GNN-AGGL-Base model. However, as will be discussed in Section 3, this level of complexity is required to tame the inherited geometric complexity of the three-dimensional setting correctly.
@@
 •  Coordinates rescaled in [1,1][-1,1].  •  Coordinates standardization:
Normalization    Mean: $\mu=0$ - Variance: $\sigma^{2}=1$
@@
Linear Layer 3 decreasing-layers of 328232-8-2 neurons. 4 decreasing-layers of 64328264-32-8-2 neurons.
@@
The training of the models has been performed in an unsupervised learning setting. The loss function that is used is the expected normalized cut, defined as:
$\mathscr{L}(Y,G)=\sum_{(i,j)\in E}\left(\dfrac{Y_{i1}(1-Y_{j1})}{\Gamma_{1}}+\dfrac{Y_{i2}(1-Y_{j2})}{\Gamma_{2}}\right),$ (1)
where $\Gamma_{k}=\sum_{i\in V}Y_{ik}D_{i}$ and $D$ is the vector of degrees of nodes. For details regarding the construction of the loss function, we refer to [27]. Concerning the construction of the dataset, we generate a training dataset of $500$ tetrahedral meshes ($125$ meshes of the unitary cube and $375$ meshes of suitable portions of them). All the meshes are generated using GMSH [53]. We adopt a split between the training set and validation set, with the training set comprising $80\%$ and the validation set comprising $20\%$ of the total dataset. As a consequence, the training set is composed of $400$ grids ($100$ meshes of the unitary cube and $300$ meshes of portions). At the same time, the validation dataset is made by the remaining $100$ grids. To guarantee good agglomeration results with our GNN models, we use a dataset with a large variety of meshes, differentiating the number of mesh elements. In the training and validation datasets, we start from coarse meshes of $600$ elements, until around 42 00042\,000 elements of mesh in the finest one. In Figure 3, we report some examples of meshes from the training set.
@@
  • Regularization, summing a $L^{2}$-regularization term to the normalized cut loss, leading to the following loss definition:
    $\hat{\mathscr{L}}(W,Y,G)=\mathscr{L}(Y,G)+\lambda\|W\|^{2}_{2}.$ (2)
@@
The training process has been performed by adopting an Adam optimizer. All the details about the training hyperparameters for the two models are reported in Table 2.
@@
Batch size 44 Samples 44 Samples
@@
Table 2: Training hyperparameters adopted for the two proposed GNN models.
[…]
@@
In Figure 4, we report the values of the loss functions for the training and validation sets for the GNN-AGGL-Base and the GNN-AGGL-Enhanced models. We can observe that using the GNN-AGGL-Enhanced model, the loss function reaches lower values on both datasets. Moreover, the training procedure is more stable compared to the two-dimensional setting [27]. This is probably a consequence of the fact that the grids of the three-dimensional training dataset are made only of tetrahedra and not of general shape.